OpenStudy (anonymous):

20b^2-62b+30=0 Solve Quadratic Equation by Factoring.

OpenStudy (anonymous):

\[20b ^{2}-62b+30\]

OpenStudy (anonymous):

just search quadratic equation calculator lol that's what helped me

OpenStudy (anonymous):

But will it show the work?

OpenStudy (dtan5457):

Well considering you don't have a quadratic calculator on a test...you can use the quadratic formula, borrow payback method. Which one sounds appealing to you..?

OpenStudy (anonymous):

Quadratic.

OpenStudy (anonymous):

Lmao! I was just recommending it damn. it literally helped me learn it but ok

OpenStudy (anonymous):

Let's solve your equation step-by-step. 20b2−62b+30=0 Step 1: Factor left side of equation. 2(5b−3)(2b−5)=0 Step 2: Set factors equal to 0. 5b−3=0 or 2b−5=0 b= 3 5 or b= 5 2 Answer: b= 3 5 or b= 5 2

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

@johnweldon1993 like a boss lol

OpenStudy (anonymous):

yep anytime this is easy to me!

OpenStudy (anonymous):

haha. I just missed too a few days at school. you are Awesome

OpenStudy (dtan5457):

@LightYagami96 What are your answers exactly..? The way you spaced it out is hard to read.

OpenStudy (anonymous):

3/5 or 5/2 is the answer

OpenStudy (dtan5457):

Right... xD

OpenStudy (dtan5457):

That's a weird way to put your fractions.